# Quantitative Comparison of Fine-Tuning Techniques for Pretrained Latent Diffusion Models in the Generation of Unseen SAR Image Concepts

### v1

@@ Abstract
This work investigates the adaptation of large pre-trained latent diffusion models to a radically new imaging domain: Synthetic Aperture Radar (SAR). While these generative models, originally trained on natural images, demonstrate impressive capabilities in text-to-image synthesis, they are not natively adapted to represent SAR data, which involves different physics, statistical distributions, and visual characteristics. Using a sizeable SAR dataset (on the order of 100,000 to 1 million images), we address the fundamental question of fine-tuning such models for this unseen modality. We explore and compare multiple fine-tuning strategies, including full model fine-tuning and parameter-efficient approaches like Low-Rank Adaptation (LoRA), focusing separately on the UNet diffusion backbone and the text encoder components. To evaluate generative quality, we combine several metrics: statistical distance from real SAR distributions, textural similarity via GLCM descriptors, and semantic alignment assessed with a CLIP model fine-tuned on SAR data. Our results show that a hybrid tuning strategy yields the best performance: full fine-tuning of the UNet is better at capturing low-level SAR-specific patterns, while LoRA-based partial tuning of the text encoder, combined with embedding learning of the <SAR> token, suffices to preserve prompt alignment. This work provides a methodical strategy for adapting foundation models to unconventional imaging modalities beyond natural image domains.
@@ 1 Introduction
Synthetic Aperture Radar (SAR) imaging is playing an increasingly central role in remote sensing, thanks to its unique physical properties and the rapidly expanding ecosystem of sensors. Unlike optical systems, SAR is an active sensor that transmits its own signal, enabling image acquisition regardless of lighting conditions. This independence from ambient illumination makes SAR particularly well-suited for continuous, all-weather monitoring. Additionally, SAR signals can penetrate certain materials such as vegetation, sand, or dry snow, giving access to information inaccessible to optical sensors. These characteristics, combined with a growing diversity of platforms, from small satellites to airborne or drone systems operating across different frequency bands, have established SAR as a key modality for various applications, including surveillance, environmental monitoring, urban mapping, and disaster assessment.
However, SAR images exhibit fundamental differences from natural optical images. First, the imaging geometry is non-conventional: SAR data is acquired in a slant-range and azimuth coordinate system, introducing distortions that depend on viewing angle and terrain relief. Second, SAR amplitude values are not governed by Gaussian distributions but instead follow speckle statistics, due to the interference phenomena intrinsic to coherent systems. This results in highly textured images and heavy-tailed statistical distributions. Moreover, the radiometric content of SAR images spans a wide dynamic range, with bright reflections (e.g., from man-made structures or double-bounce effects) coexisting with very low-return regions (e.g., calm water). These properties are further modulated by acquisition parameters such as polarization, incidence angle, resolution, and wavelength, making SAR data complex to model.
This diversity and variability challenge conventional computer vision methods, particularly those developed and trained on optical imagery. One promising direction lies in the simulation of SAR data, which enables algorithm testing, rare scenario generation, and dataset augmentation. Recent advances in generative models, especially diffusion-based approaches, have revived interest in image synthesis. Diffusion models have been shown to surpass GANs in realism and diversity.
Among these, latent diffusion models represent a reference class of generative models. Instead of performing the denoising process directly in pixel space — which is computationally intensive and less efficient for high-resolution data — they work in a compressed space called the latent space. This latent space is learned through a Variational Autoencoder (VAE), whose encoder maps images into lower-dimensional representations, and whose decoder reconstructs images from this space. The generative process itself is handled by a UNet-based denoising network, trained to reverse a noise corruption process (called forward diffusion process) in the latent domain. This reverse diffusion process is conditioned on text embeddings produced by text encoders, which allows the model to associate semantic meaning with the denoising trajectory. These models, when pretrained on massive optical datasets, become powerful generative priors, capable of synthesizing photorealistic scenes and semantically coherent layouts. However, they lack any native knowledge of SAR-specific image structures, noise patterns, or physical signal distributions.
In this work, we explore the adaptation of such a pre-trained latent diffusion model to the SAR modality. Starting from a domain-specific dataset of approximately 100,000 high-resolution airborne SAR images, acquired with ONERA’s SETHI sensor, we aim to investigate how such a model can be fine-tuned to synthesize realistic radar imagery from textual prompts. The scale of this dataset places us in an intermediate regime: substantially larger than typical few-shot setups, yet considerably smaller than the hundreds of millions of samples used for pretraining foundation models, making it a relevant test case for transfer learning strategies.
This brings us to a central question: what fine-tuning strategy should be employed to effectively adapt a pre-trained model to such an unfamiliar modality? Should all components of the model be retrained, or can we rely on more lightweight approaches such as Low-Rank Adaptation (LoRA) to update only a subset of parameters? To answer this, we carry out a comparative study of several fine-tuning strategies, applied selectively to the text encoder and the UNet diffusion backbone.
@@
An equally important challenge is how to evaluate the quality of generated SAR images. Traditional visual metrics are ill-suited to this task, as they assume natural image statistics. In response, we propose a new evaluation framework, combining statistical distribution comparisons, texture analysis via Gray-Level Co-occurrence Matrices, and semantic alignment using a CLIP model fine-tuned on SAR-caption pairs.
@@
The remainder of the paper is organized as follows. We begin by reviewing related work in generative modeling and fine-tuning approaches, with a focus on parameter-efficient adaptation methods. We then describe the construction of our SAR dataset and the pre-processing pipeline applied to the SETHI archive. The methodological section details the diffusion model architecture, the fine-tuning strategies considered, and our proposed evaluation metrics. This is followed by a presentation of our experimental results, including both quantitative analyses and visual examples of generated SAR images. We conclude with a discussion of our findings, limitations, and perspectives for future work.
@@ 2 Related work
Recent years have seen the emergence of large-scale generative models capable of synthesizing images from natural language prompts. These models, commonly referred to as foundation models, are typically pre-trained on massive datasets and designed to capture high-level semantic alignment between visual and textual modalities. Their success has led to interest in adapting them to specialized domains, such as medical imaging or remote sensing. However, extending these models to unconventional modalities such as SAR, which differs both structurally and statistically from natural images, remains a largely underexplored challenge. In this section, we review the relevant efforts in basic vision language modeling and the fine-tuning strategies developed to adapt them effectively.
@@ 2.1 Foundation Generative Vision-Language Models
Vision-Language Models (VLMs) have demonstrated strong capabilities in learning joint embeddings and generative alignments across visual and textual modalities. Early models such as CoCa [1] combine contrastive and generative objectives to jointly align and synthesize, while more recent architectures like CM3Leon and Chameleon [2] implement early fusion designs, allowing unified multimodal generation through transformer-based architectures. These systems can generate both image and text outputs conditioned on joint multimodal inputs.
In parallel, a distinct family of models specifically focuses on text-to-image generation. Within this space, Stable Diffusion [3], Flux [4], Imagen [5], and Parti [6] represent several lines of research exploring different generation mechanisms: respectively, latent-space diffusion, pixel-space diffusion, and autoregressive modeling. Among them, latent diffusion models such as Stable Diffusion have gained attention for their ability to generate high-resolution images with lower computational cost. This is achieved by operating in a compressed latent space, learned via a Variational Autoencoder (VAE), rather than directly in pixel space.
Stable Diffusion XL (SDXL) [7], in particular, is a flexible, open-source latent diffusion model consisting of three main components: (i) a VAE that compresses images into latent representations, (ii) a dual text encoder pipeline to transform prompts into embeddings, and (iii) a UNet backbone that performs conditional denoising in the latent domain. Numerous extensions have been proposed to control its generation process, including spatial guidance methods such as ControlNet [8], and global image-based conditioning methods like IP-Adapter [9]. However, current research has mostly focused on natural image domains, and little is known about the model’s ability to learn physically grounded or domain-specific concepts such as those present in SAR imagery.
@@ 2.2 Fine-tuning approaches
As the size of pre-trained Vision-Language Models (VLMs) continues to grow, full fine-tuning— i.e., updating all model parameters — becomes increasingly impractical due to memory, compute, and data constraints. To address this, a family of Parameter-Efficient Fine-Tuning (PEFT) techniques has emerged, aiming to adapt large-scale models by updating only a small fraction of their weights parameters.
@@
One of the most widely used PEFT strategies is Low-Rank Adaptation (LoRA) [10], which injects trainable low-rank matrices into the linear layers of the model. LoRA enables effective adaptation while keeping the majority of weights frozen, drastically reducing memory usage. Extensions such as QLoRA [11] and DoRA [12] further optimize efficiency by combining low-rank decomposition with quantization or weight reparameterization, especially for large language models (LLMs), and are increasingly being explored in vision and multimodal settings.
@@
Other approaches focus on prompt-level conditioning rather than internal parameter modification. Prompt-based tuning methods such as CoOp [13] and VPT [14] learn input embeddings or prompts that guide the model without altering its architecture. These techniques are lightweight and adaptable, but may can’tb represent entirely new visual domains when semantic gaps are large.
@@
In contrast, DreamBooth [15] enables explicit concept injection by associating new visual identities or styles with custom textual tokens. This method has been successfully applied to generate highly specific outputs from small datasets — for instance, [16] fine-tuned Stable Diffusion 3 on 300 samples of Jamini Roy-style paintings, achieving culturally accurate synthesis. Further control can be obtained using ControlNet and IPAdapter, which allow generation to be conditioned on structural priors such as edge maps, segmentation, or depth — especially effective for layout-sensitive domains.
@@
Another promising method is textual inversion [17], which learns new visual concepts directly in the embedding space of the text encoder, without modifying the image generator. When combined with DreamBooth and LoRA, as in [18], it enables joint control over modality and identity — for example, generating paired visible-infrared images from shared prompts like "a [modality] photo of a [person] person".
@@
Collectively, these techniques have made it feasible to adapt powerful diffusion models such as Stable Diffusion to novel visual concepts, ranging from new objects to artistic styles, while requiring relatively modest amounts of data and compute. However, most existing applications remain confined to natural image domains, and focus on concept categories that are semantically close to those seen during pre-training (e.g., human faces, animals, or art styles). The adaptation of foundation models to structurally and statistically distinct modalities — such as depth images, medical scans, or Synthetic Aperture Radar (SAR)—remains a largely unexplored and technically challenging frontier.
@@ 2.3 Generative Foundation models in Remote Sensing
In the field of remote sensing, most recent generative Vision-Language Models (VLMs) have focused primarily on optical imagery, with limited or no support for Synthetic Aperture Radar (SAR) data. Several foundation models have been trained from scratch on large-scale optical satellite image datasets, including RS5M and GeoRSCLIP [19], DiffusionSat [20], MetaEarth [21], CRS-Diff [22], and HSIGene [23], the latter targeting hyperspectral image generation. While most of these models are trained for representation learning, zero-shot classification, or retrieval in the optical domain, only a few are designed for image generation — and even fewer extend to radar-based modalities such as Synthetic Aperture Radar (SAR).
To our knowledge, Text2Earth [24] is the first foundation model that incorporates both SAR and optical data for text-to-image generation. However, its SAR component relies on synthetic radar - like images produced via Pix2Pix translation from RGB inputs, rather than using real SAR measurements that include speckle noise, geometric distortions, and backscatter - specific statistical properties. As such, the model does not capture the full complexity of radar signal characteristics.
Other models, such as SARChat-InternVL2.5-8B [25], have focused on improving multimodal understanding of SAR imagery through conversational tasks like description, counting, or spatial reasoning. However, this model is not able to do image generation, and its training data is limited to open-source object detection benchmarks. It does not address large-scale SAR image synthesis nor generalization across acquisition conditions.
In general, the development of generative foundation models for SAR remains in its infancy, especially for high-resolution data and for tasks beyond classification or detection. While some efforts have explored training from scratch, such approaches are computationally prohibitive and require extensive domain - specific data. In contrast, adapting pre-trained generative models - originally trained on optical images - to the SAR domain via fine-tuning offers a more scalable and practical alternative. Yet, this path remains largely underexplored. Our work positions itself in this gap, investigating how such pre-trained models can be effectively adapted to synthesize realistic SAR imagery guided by textual prompts.

### v2

@@ Abstract
We present a framework for adapting a large pretrained latent diffusion model to high-resolution Synthetic Aperture Radar (SAR) image generation. The approach enables controllable synthesis and the creation of rare or out-of-distribution scenes beyond the training set. Rather than training a task-specific small model from scratch, we adapt an open-source text-to-image foundation model to the SAR modality, using its semantic prior to align prompts with SAR imaging physics (side-looking geometry, slant-range projection, and coherent speckle with heavy-tailed statistics). Using a 100k-image SAR dataset, we compare full fine-tuning and parameter-efficient Low-Rank Adaptation (LoRA) across the UNet diffusion backbone, the Variational Autoencoder (VAE), and the text encoders. Evaluation combines (i) statistical distances to real SAR amplitude distributions, (ii) textural similarity via Gray-Level Co-occurrence Matrix (GLCM) descriptors, and (iii) semantic alignment using a SAR-specialized CLIP model. Our results show that a hybrid strategy—full UNet tuning with LoRA on the text encoders and a learned token embedding—best preserves SAR geometry and texture while maintaining prompt fidelity. The framework supports text-based control and multimodal conditioning (e.g., segmentation maps, TerraSAR-X, or optical guidance), opening new paths for large-scale SAR scene data augmentation and unseen scenario simulation in Earth observation.
@@ 1 Introduction
Synthetic Aperture Radar (SAR) has become a key modality in Earth observation, no longer restricted to governmental or defense institutions. It is supported by a growing diversity of platforms—from small satellites to airborne and drone systems operating across multiple frequency bands—and provides day-night, all-weather imaging for environmental monitoring, urban mapping, surveillance, and disaster assessment.
In this context, synthetic data are crucial for model testing, algorithm and sensor development, operational deployment, and data interpretation. Due to the variety of SAR applications, different simulation approaches are needed. In the literature, physics-based simulators typically require extensive information about objects’ geometry and material, radar parameters, and other specific details, which are often difficult to obtain, for example, RaySAR from Auer et al. (2016), SARCASTIC v2.0 from Woollard et al. (2022), and MOCEM COCHIN et al. (2008). Most existing tools focus on the simulation of isolated targets or objects for detection or classification tasks. For large scene SAR simulation, a promising direction is the use of generative AI, enabling scalable labeled dataset augmentation with both controllable image synthesis and the creation of unseen, rare or challenging scenarios beyond existing datasets.
While Generative Adversarial Networks (GANs) have been explored mainly for isolated-target SAR image generation and augmentation (e.g., ships, vehicles, equipment models), they largely reproduce low-level statistics and offer limited multimodal control, constraining scene-level composition Zou et al. (2020), Gao et al. (2024), Liu et al. (2018). In contrast, Latent diffusion models (LDMs) presented in Rombach et al. (2022) learn to invert a noise process in a compact latent space, enabling efficient generation with self- and cross-attention conditioning. As of foundation LDM, Stable Diffusion by Rombach et al. (2022) uses a Variational Autoencoder (VAE) to map images to latent representations; text encoders provide embeddings that condition an attention-based UNet during denoising. Pretrained on large web-sourced text–image pairs of optical images, Stable Diffusion is a powerful generative model capable of synthesizing photorealistic scenes and layouts.
Adapting such foundation models to SAR remains challenging because SAR imagery differs fundamentally from optical training datasets. SAR is acquired in a side-looking geometry with slant-range and azimuth coordinates (range–Doppler), producing layover, foreshortening, and shadow effects that depend on incidence angle and terrain. Coherent imaging creates speckle and heavy-tailed amplitude statistics, while the radiometric dynamic range is wide, with bright man-made backscatter coexisting with very low-return areas (e.g., calm water). These properties vary with polarization, incidence, resolution, and wavelength, complicating direct transfer and motivating domain adaptation.
Thus, we present an adaptable framework to fine-tune an open-source pretrained LDM—specifically Stable Diffusion XL (SDXL) by Podell et al. (2023)—to the SAR modality. Our goal is to preserve the semantic prior and compositional abilities of the base text-to-image model while aligning generation with SAR imaging physics. To reduce acquisition-dependent variability, we curate a consistent high-resolution dataset ( 40 cm, X-band) with similar incidence angles and processing, resulting in  100,000 airborne SAR images acquired with ONERA’s SETHI sensor Baqué et al. (2019).
To make synthetic SAR data useful for downstream tasks, the generative model must go beyond the training set by composing novel scene configurations on demand. We use the pretrained model’s semantic prior to describe and compose diverse environments (e.g., urban, forested, coastal) with spatial relations (e.g., “along,” “near,” “to the left of”). Our adaptation preserves the model’s language and relational understanding while aligning generation with SAR imagery, ensuring outputs are physically plausible SAR scenes preserving rather than synthesizing grayscale optical renderings with added speckle.
To find a non-trivial balance between overtraining—which would ensure a good understanding of SAR physics but would lead to a loss of model plasticity—and undertraining—which would result in a simple stylization of optical images—it is necessary to study the fine-tuning process carefully to identify the right trade-off. To address this, we investigate various fine-tuning approaches on the UNet backbone, VAE, and Text Encoders of Stable Diffusion XL—including adjustments to hyperparameters, weight updates, embedding learning (with a SAR-specific token), and loss regularization. These strategies aim to learn SAR-specific representations that capture speckle, texture, and reflectivity coherence.
@@
Finally, an equally important challenge is how to evaluate the quality of generated SAR images. Traditional visual metrics in AI are ill-suited to this task, as they assume natural image statistics. In response, we propose a new evaluation framework, combining statistical distribution comparisons, texture analysis via Gray-Level Co-occurrence Matrices, and semantic alignment using a CLIP model fine-tuned on SAR-caption pairs. Using these metrics, we compare training configurations to make model behavior more explainable and to use the framework to other latent diffusion models (LDMs).
This adaptation offers numerous practical applications, including the generation of unseen, rare, and challenging scenes beyond existing datasets. It also enables controllable image synthesis with image-to-image generation while preserving spatial or statistical priors, improving tasks like adding structured spatial details and refining physics-based simulations. Here, we show that our method improves the realism of synthetic SAR imagery generated by ONERA’s EMPRISE simulator and of synthesis conditioned on TerraSAR-X imagery.
@@
The paper first presents related work in generative modeling and fine-tuning methods, focusing on parameter-efficient adaptation. After describing the training dataset, the methodology section outlines the diffusion model architecture, fine-tuning strategies, and evaluation metrics. Experimental results, including quantitative analysis and visual examples of generated SAR images, are presented next. Our work on improving the realism of synthetic SAR images generated by ONERA’s EMPRISE simulator and enhancing TerraSAR-X satellite acquisitions is then discussed. It concludes with a discussion of our findings, limitations, and future directions.
@@ 2 Related work
Recent years have seen the emergence of large-scale generative models capable of synthesizing images from natural language prompts. These models, commonly referred to as foundation models, are typically pretrained on massive datasets and designed to capture high-level semantic alignment between visual and textual modalities. Their success has led to interest in adapting them to specialized domains, such as medical or optical remote sensing imaging. However, extending these models to unconventional modalities such as SAR, which differs both structurally and statistically from natural images, remains a largely underexplored challenge. In this section, we review the relevant efforts in basic vision language modeling and the fine-tuning strategies developed to adapt them effectively.
@@ 2.1 Foundation Generative Vision-Language Models
Vision-Language Models (VLMs) have demonstrated strong capabilities in learning joint embeddings and generative alignments across visual and textual modalities. Early models such as CoCa by Yu et al. (2022) combine contrastive and generative objectives to jointly align and synthesize, while more recent architectures like CM3Leon and Chameleon by Team (2025) implement early fusion designs, allowing unified multimodal generation through transformer-based architectures. These systems can generate both image and text outputs conditioned on joint multimodal inputs.
In parallel, a distinct family of models specifically focuses on text-to-image generation. Within this category, Stable Diffusion by Rombach et al. (2022), Flux by Labs (2024), Imagen by Saharia et al. (2022), and Parti by Yu et al. (2023) represent several lines of research exploring different generation mechanisms: respectively, latent-space diffusion, pixel-space diffusion, and auto-regressive modeling. Among them, latent diffusion models such as Stable Diffusion have gained attention for their ability to generate high-resolution images with lower computational cost. This is achieved by operating in a compressed latent space, learned via a Variational Autoencoder (VAE), rather than directly in pixel space.
Stable Diffusion XL (SDXL) by Podell et al. (2023), in particular, is a flexible, open-source latent diffusion model consisting of three main components: (i) a VAE that compresses images into latent representations, (ii) a dual text encoder pipeline to transform prompts into embeddings, and (iii) a UNet backbone that performs conditional denoising in the latent domain. Numerous extensions have been proposed to control its generation process, including spatial guidance methods such as ControlNet by Zhang et al. (2023), and global image-based conditioning methods like IP-Adapter by Ye et al. (2023). However, current research has mostly focused on optical image domains, and little is known about the model’s ability to learn physically grounded or domain-specific concepts such as those present in SAR imagery.
@@ 2.2 Fine-tuning approaches
As the size of pretrained Vision-Language Models (VLMs) continues to grow, full fine-tuning—i.e., updating all model parameters—becomes increasingly impractical due to memory, compute, and data constraints. To address this, a family of Parameter-Efficient Fine-Tuning (PEFT) techniques has emerged, aiming to adapt large-scale models by updating only a small fraction of their weights parameters.
@@
One of the most widely used PEFT strategies is Low-Rank Adaptation (LoRA) presented by Hu et al. (2021), which injects trainable low-rank matrices into the linear layers of the model. LoRA enables effective adaptation while keeping the majority of weights frozen, drastically reducing memory usage. Extensions such as QLoRA by Dettmers et al. (2023) and DoRA by Wang et al. (2024) further optimize efficiency by combining low-rank decomposition with quantization or weight reparameterization, especially for large language models (LLMs), and are increasingly being explored in vision and multimodal settings.
@@
Other approaches focus on prompt-level conditioning rather than internal parameter modification. Prompt-based tuning methods such as CoOp by Zhou et al. (2022) and VPT by Jia et al. (2022) learn input embeddings or prompts that guide the model without altering its architecture. These techniques are lightweight and adaptable, but may not represent entirely new visual domains when semantic gaps are large.
@@
In contrast, DreamBooth presented by Ruiz et al. (2023) enables explicit concept injection by associating new visual identities or styles with custom textual tokens. This method has been successfully applied to generate specific outputs from small datasets — for instance, Agrawal and Banerjee (2025) fine-tuned Stable Diffusion 3 on 300 samples of Jamini Roy-style paintings, achieving culturally accurate synthesis. Further control can be obtained using ControlNet and IPAdapter, which allow generation to be conditioned on structural priors such as edge maps, segmentation, or depth — especially effective for layout-sensitive domains.
@@
Another promising method is textual inversion presented by Gal et al. (2022), which learns new visual concepts directly in the embedding space of the text encoder, without modifying the image generator. When combined with DreamBooth and LoRA, as in the paper from Dai et al. (2025), it enables joint control over modality and identity — for example, generating paired visible-infrared images from shared prompts like "a [modality] photo of a [person] person".
@@
Collectively, these techniques have made it feasible to adapt powerful diffusion models such as Stable Diffusion to novel visual concepts, ranging from new objects to artistic styles, while requiring relatively modest amounts of data and compute. However, most existing applications remain confined to optical image domain and focus on concept categories that are semantically close to those seen during pre-training (e.g., human faces, animals, or art styles). But these methods are not well-suited for learning entirely new domains, such as SAR imagery, which is structurally and statistically distinct. Indeed, LoRA-based method alone can’t capture the spatial complexities of SAR data, including Rayleigh noise and imaging geometry. Prompt-based tuning methods like CoOp and VPT are limited when dealing with large semantic gaps, such as those between optical and SAR domains. DreamBooth or Textual inversion, though effective for small datasets, struggle with SAR’s features, requiring larger, high-resolution data.
@@ 2.3 Generative Foundation models in Remote Sensing
In the field of remote sensing, most recent generative Vision-Language Models (VLMs) have focused primarily on optical imagery, with limited or no support for Synthetic Aperture Radar (SAR) data. Several foundation models have been trained from scratch on large-scale optical satellite image datasets, including RS5M and GeoRSCLIP by Zhang et al. (2024), DiffusionSat by Khanna et al. (2024), MetaEarth by Yu et al. (2024), CRS-Diff by Tang et al. (2024), and HSIGene by Pang et al. (2024), the latter targeting hyperspectral image generation. While most of these models are trained for representation learning, zero-shot classification, or retrieval in the optical domain, only a few are designed for image generation — and even fewer extend to radar-based modalities such as Synthetic Aperture Radar (SAR).
To our knowledge, Text2Earth by Liu et al. (2025) is the first foundation model that incorporates both SAR and optical data for text-to-image generation. However, its SAR component relies on synthetic radar-like images produced via Pix2Pix translation from RGB inputs, rather than using real SAR measurements that include speckle noise, geometric distortions, and backscatter-specific statistical properties. As such, the model does not capture the full complexity of radar signal characteristics.
Other models, such as SARChat-InternVL2.5-8B presented by Ma et al. (2025), have focused on improving multimodal understanding of SAR imagery through conversational tasks like description, counting, or spatial reasoning. However, this model is not able to do image generation, and its training data is limited to open-source object detection benchmarks. It does not address large-scale SAR image synthesis nor generalization across acquisition conditions.
In general, the development of generative models for SAR is still in its early stages, particularly for high-resolution data. While some efforts have explored training from scratch, such approaches are computationally prohibitive and require extensive domain-specific data. In contrast, adapting pretrained generative models - originally trained on optical images - to the SAR domain via fine-tuning offers a more scalable and practical alternative. Yet, this path remains largely underexplored. Our work positions itself in this gap, investigating how such pretrained models can be effectively adapted to synthesize realistic SAR imagery guided by textual prompts.